OpenStudy (anonymous):

prove that for all n>0, n!>=2^(n-1) using induction

OpenStudy (anonymous):

basis step using n=1 is true

OpenStudy (anonymous):

plugging in n+1

OpenStudy (anonymous):

\[(n+1)n!\ge2^{n-1}(2)\]

OpenStudy (anonymous):

\[n!\ge2^{n-1}(\frac{ 2 }{ n+1 })\]

OpenStudy (anonymous):

we know n>0, n+1>1, n+1>=2 (because n cannot be non-integer?)

OpenStudy (anonymous):

here's where I get stuck

hartnn (hartnn):

you assume the statement true for n so, you have n! > = 2^(n-1) now you need to prove (n+1)! > = 2^n [replacing n by n+1] right ? (n+1)n! > = 2^(n-1) (n+1) (n+1)! > = 2^n [(n+1)/2] since , n > 0 (n+1) /2 > 0 so,(n+1)! > = 2^n

OpenStudy (anonymous):

ok, in the second line, is n+1 in the exponent on RHS

hartnn (hartnn):

no... n+1 was multiplied on both sides...its not in exponents.

OpenStudy (anonymous):

don't we replace all instances of n with n+1?

OpenStudy (anonymous):

doing it on paper the way you suggested, i think i see what you mean, multiplying by n+1 on both sides is the way of proving for n+1 in inductive step

OpenStudy (anonymous):

i hope u understand my solution, let me know if u need help

OpenStudy (anonymous):

tushara, do you agree with the other answer in this question?

OpenStudy (anonymous):

the only thing m not sure about with the other solution that hartn did is that,,,, (n+1) was multiplied with the power.... i thot u cant do that, u can multiply it with the whole eqn on the rhs side but u cant multiply it just to the power

OpenStudy (anonymous):

in my solution i wrote k+1=k=n.... scratch dat, it shud be....k+1=n+1 and from there it follows right

OpenStudy (anonymous):

are we allowed to draw the conclusion that the RHS changes similarly to the LHS, where RHS must be 2^(k+1-1)

OpenStudy (anonymous):

yes , ive seen other proofs which were done dis way

OpenStudy (anonymous):

how do you end up with k - k - 1 for k!

OpenStudy (anonymous):

is it because there are k items?

OpenStudy (anonymous):

because wen u do the factorial... u end up multiplying my 1 in the end right? so k-(k-1) will give 1..... so say k was 3.... 3-(3-1)=1 so the last number u multiply by is 1

OpenStudy (anonymous):

its important to remember how to expand a factorial... its always: n! = n*(n-1)*(n-2)*...........n-(n-1)

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

is there a technical term for proving in the manner you did, "take notice"

OpenStudy (anonymous):

nice... HI5!

OpenStudy (anonymous):

wat is the technical term?

OpenStudy (anonymous):

i'm asking you...my professor may request my source

OpenStudy (anonymous):

oh uhm sorry, i read ur reply wrong...... uhm no... i cant think of putting that in a more proffesional way.... u shud ask ur friends how dey went about it

OpenStudy (anonymous):

will do, i'll call it something like "rule of equalities"

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

thanks again

hartnn (hartnn):

\( n! > = 2^{(n-1)}\\(n+1)n! > = 2^{(n-1)} (n+1) \\ (n+1)! > = 2^n [(n+1)/2]\) never multiplied anything with exponent...

OpenStudy (anonymous):

your proof seems more conventional

OpenStudy (anonymous):

can you tell me how you came to the conclusion since , n > 0 (n+1) /2 > 0 so,(n+1)! > = 2^n

hartnn (hartnn):

n > 0 (n+1) /2 > 0 is pretty logical. if n is positive no., n/2 +1/2 will also be positive .

OpenStudy (anonymous):

yes, got the logical part

hartnn (hartnn):

(n+1)!>=2^n[(n+1)/2] its like x>= ab if b is positive and >=1, then x will be >= a also. so, we need (n+1)/2 > =1 since n is atleast =1, (n+1) is atleast 2 and (n+1)/2 is atleast 1 so, (n+1)/2 >=1

OpenStudy (anonymous):

how does (n+1)! > = 2^n relate to our inductive assumption?

hartnn (hartnn):

3 steps for induction. 1) prove it for n=1 2) assume it for n=n 3) using 2) prove it for n=n+1 ...some people use new variable 'k' to avoid confusion....

OpenStudy (anonymous):

oh, i must be blind, 2^n === 2^(n+1-1)

OpenStudy (anonymous):

it's proved then for n+1

hartnn (hartnn):

yes.

OpenStudy (anonymous):

one more qustion, we know that (n+1)/2>0, but what if it is a very large number? could htat make the RHS bigger than LHS?

OpenStudy (anonymous):

(n+1)! >= 2^n (1,000,000,000) for example?

hartnn (hartnn):

RHS =0 any large +ve no. +1/2 = +ve no > 0 so RHS can never be > LHS

OpenStudy (anonymous):

ok, thanks

hartnn (hartnn):

or was your doubt something else ?

OpenStudy (anonymous):

no, it is reasonable to think that LHS will grow larger than RHS, if the n becomes larger